OpenStudy (anonymous):

y'=ycosx/(1+y^2), y(0)=1 Here is my method. y=integral of ycosx/(1+y^2) dx 1+y^2=sinx+C Why is this wrong?

ganeshie8 (ganeshie8):

Look at ur right hand side : y'=ycosx/(1+y^2)

ganeshie8 (ganeshie8):

when u dont knw "y", how can u take integral of "y" with.respect.to. ANOTHER VARIABLE x ?

ganeshie8 (ganeshie8):

your task is to find "y", so first step is to write y' as dy/dx

ganeshie8 (ganeshie8):

\(\large y'=\frac{y\cos x}{1+y^2} \) \(\large \frac{dy}{dx}=\frac{y\cos x}{1+y^2} \)

ganeshie8 (ganeshie8):

separate variables by cross multiplying

ganeshie8 (ganeshie8):

\(\large \frac{1+y^2}{y}~dy = \cos x ~dx\) \(\large \frac{1}{y} + y ~dy = \cos x ~dx\)

ganeshie8 (ganeshie8):

Integrate both sides now

ganeshie8 (ganeshie8):

\(\large \int \frac{1}{y} + y ~dy = \int \cos x ~dx\) \(\large \ln y + \frac{y^2}{2} = \sin x + C\)

ganeshie8 (ganeshie8):

see if that makes more or less sense.. .